OpenStudy (anonymous):

Write the quadratic function in vertex form. y = x2 + 4x + 7

OpenStudy (amistre64):

complete the square

OpenStudy (amistre64):

or for a more general approach: when x = -b/(2a) y = a(-b/2a)^2 +b(-b/2a) +c = b^2a/(4a^2) - b^2/(2a) +c = b^2/(4a) - b^2/(2a) +c = b^2/(4a) - 2b^2/(4a) + 4ac/(4a) = (-b^2 + 4ac)/(4a)

OpenStudy (amistre64):

y can be simplified even further as y = -b^2/(4a) + 4ac/(4a) y = c - b^2/(4a)

OpenStudy (anonymous):

add and subtract 4 y=x^2+4x+4-4+7 since \[(a+b)^2=a^2+b^2+2ab\] y=(x+2)^2+3 y-3=(x+2)^2